# Rule-driven News Captioning

### 11 Mar 2024 v2

@@ Abstract
News Image Captioning, Rule-Driven, Large-Scale Pre-trained Model
@@ I Introduction
To tackle this problem, we propose the rule-driven news captioning method, which can generate image descriptions following designated rule signal. Our method consists of three components: 1) Named Entity Extraction. We use the pre-trained named entity recognition (NER) technique to extract named entities from news articles. Then, the CLIP model is employed to compute the image-text similarity to match the named entities that closely align with the news image content. 2) News Rule Construction. We design the news-aware semantic rule for the descriptions. This rule incorporates the primary action depicted in the image (e.g., “performing”) and the set of roles played by named entities surrounding the action (e.g., “Agent” and “Place”). Meanwhile, it can present a clear affiliation between generic objects (e.g., “man”) in images and named entities (e.g., “Pedro Sa Moraes”) in news articles. 3) Caption Generation. We incorporate the news-aware semantic rule into the large-scale pre-trained model, BART, using the prefix-tuning strategy. Specifically, BART can be used as the pre-trained multi-modal summarization model [7] to generate captions for both images and news articles. We embed the semantic rule into BART’s multiple encoder layers. As a result, BART is effectively guided to generate news sentences that adhere to the rules of news reporting.
We conduct comprehensive experiments on two publicly available news datasets, i.e., GoodNews [5] and NYTimes800k [3]. Experimental results show that our method can achieve competitive or state-of-the-art performance against the existing methods on all benchmarks. Especially under CIDEr metric, our method can achieve significant improvement over the best competitor [6] by 8.93% on GoodNews. Ablation studies further examine the efficacy of each component. Besides, we visualize the generated news-aware semantic rules, to clearly observe the named entities’ changes improved by our method. In summary, we have the following main contributions:
@@ II-A General Image Captioning
The image captioning task is a challenging task that combines computer vision and natural language processing technology [8], [9], [10]. Vinyals et al. [8] first use CNN to encode the image, and then use LSTM to generate the corresponding image description. Yao et al. [11] propose GCN [12][13] plus LSTM. Li et al. [9] adopt a framework based on CNN-Transformer. They introduce EnTangled Attention in Transformer to integrate visual and semantic information more effectively. Compared with the CNN-Transformer framework, Liu et al. [14] achieve more efficient global context modeling through a full Transformer model. Wang et al. [15] propose CrowdCaption, a crowd scenes caption dataset, and devise a method for CrowdCaption that can generate crowd-specific detailed descriptions. Zhang et al. [16] introduce the Adaptive-Attention module in the Transformer decoder to measure the contribution of visual cues and language cues to caption generation. Wu et al. [1] use segmentation features as additional information sources to enhance the contribution of visual information in the prediction process. Cornia et al. [17] design Universal Captioner, which adopts network-level noise data for training and uses the long-tail concept to describe images. Jing et al. [18] utilize the implicit external knowledge contained in the dataset to generate video captions. Although traditional image captioning task performs very well, this method can only generate general descriptions and lacks specific contextual information. In comparison, our method uses image-article pairs as input, and with the help of the background knowledge contained in the news articles, it can generate descriptions containing named entities, which enhances the practicality and accuracy of the descriptions.
@@ II-B News Image Captioning
In recent years, news image captioning task has received increasing attention. Compared with the traditional image captioning task, the news image captioning task can effectively combine the image content and the background knowledge of the news article to generate captions containing named entities. Early research work often generates captions in two steps [5], [19]. Firstly, a caption with placeholders is generated, then the correct named entities are selected from news articles to replace the placeholders in the caption. For example, Biten et al. [5] use different types of named entities as placeholders to generate the template caption, e.g., PER performs in GPE on DATE. Then they select named entities from the sentences with the highest attention values in the news articles to fill these placeholders.
Recently, an increasing number of works have adopted end-to-end models. Hu et al. [20] implement a text information aggregation approach from the sentence-level to the word-level to generate entity-aware captions. Tran et al. [3] utilize RoBERTa [21] and ResNet-152 [22] to encode news articles and images respectively, and handle uncommon words by using byte-pair encoding in the Transformer decoder. Zhao et al. [23] propose to mine external web knowledge to construct a multi-modal knowledge graph, thereby associating visual objects with named entities. Yang et al. [4] design a set of template components based on the six elements of news (i.e., who, when, where, what, why, how) to guide the model in generating news captions that comply with journalistic guidelines. Zhang et al. [6] utilize a multi-modal entity prompting mechanism to fine-tune the large-scale pre-trained model. Zhou et al. [24] generate accurate captions by combining key local context containing important named entities with global context that can describe the entire article.
@@
Figure 2: Overview of the proposed method. (a) Named Entity Extraction. We use the Named Entity Recognition (NER) model to extract named entities and their categories from news articles. Then, the CLIP model is applied to calculate the image-text similarity score, filtering the named entities accordingly. (b) News Rule Construction. We design the news-aware semantic rule, which includes the primary action depicted in the image (e.g., “performing”) and the named entities involved in the action (e.g., “Ms. Micucci”, “Ms. Lindhome” and “theater”) along with their corresponding roles (e.g., “Agent” and “Stage”). (c) Caption Generation. We integrate the news-aware semantic rule, news article, and image, into BART, a large-scale pre-trained model, using the prefix-tuning strategy, for the news caption generation.
@@ II-C Large-Scale Pre-trained Models
Large-scale pre-trained models have received widespread attention [25], [26], [27], [28], [29], [30], [7]. These models are usually pre-trained on large-scale datasets, and then fine-tuned for specific tasks to better adapt to various complex application scenarios. They have demonstrated excellent results in tasks such as common sense reasoning [25], [26] and visual question answering [31], [32], [33]. Among them, CLIP [30] leverages 400 million image-text pairs gathered from the internet, and uses contrastive learning [34], [35] to achieve cross-modal semantic alignment between image features and text features. BART [7] is a large-scale pre-trained language model based on Transformer. Our approach enables news description generation by injecting news-aware semantic rule into multiple encoder layers of BART, using the prefix-tuning strategy.
@@ III-A Problem Formulation
In order to generate captions that adhere to fundamental rules of news reporting, we first design the news-aware semantic rule, which is filled using images and news articles. Then, we integrate this semantic rule into the large-scale pre-trained model (e.g., BART), where multiple encoder layers are embedded with news-aware semantic rule, to generate image descriptions. Formally, Eq. 1 is re-written as:
@@ III-B Named Entity Extraction
Named entities play a crucial role in enhancing the accuracy and richness of news captions by providing specific person’s name (PER), organization (ORG), and location (LOC). In this section, we extract named entities and their corresponding types from given news articles, which are then used in subsequent modules. We first employ a pre-trained named entity recognition (NER) model  [^1] https://huggingface.co/Davlan/distilbert-base-multilingual-cased-ner-hrl to extract a set of entities $E=\{e_{1},e_{2},\dots,e_{M}\}$ from the article $A$, where $M$ is the number of named entities. As shown in Fig. 2(a), to match the named entities that closely align with the news image content, we use the CLIP model [30] to compute the similarity scores between the image $I$ and each entity $e_{i}\in E$, and then select the top-$k$ entities to build the new entity set $E_{k}$:
$E_{k}=\{e_{i}\}_{i=1}^{k}=\mathrm{arg~{}Top\mbox{-}k}(\mathrm{CLIP}(I,E))$ (3)
where $\mathrm{CLIP}(I,E)$ aims to calculate image-text matching scores by the pre-trained CLIP model [30].
@@ III-D Caption Generation
We employ the frozen CLIP model to encode the image $I$ and the semantic rule $F$ independently, and obtain the visual features and the rule features, respectively. The former is embedded into the visual vector $v$ by a multi-layer perceptron (MLP), while the latter is transformed into the rule vector $f$ through a linear layer:
$v=\mathrm{MLP}(\mathrm{CLIP_{image}}(I))$ (6)
$f=\mathrm{Linear}(\mathrm{CLIP_{text}}(F))$ (7)
where $\mathrm{CLIP_{image}}(\cdot)$ is the image encoder; $\mathrm{CLIP_{text}}(\cdot)$ is the text encoder. For the article representation, we use the tokenizer to convert the news article into a collection of tokens $X^{A}=\{x_{1},x_{2},\dots,x_{N}\}$, where $x_{i}$ represents the $i$-th token in the article. We feed $X^{A}$ into an embedding layer and obtain the article vector $x^{A}$.
Essentially, the news captioning task requires adhering to some fundamental rules of news reporting, which is significantly different from the previous BART objective for the general summarization task. In this section, we utilize the prefix-tuning strategy [40]. This lightweight fine-tuning strategy involves optimizing a tiny task-specific vector called “prefix”, which can inject the task-specific structural information during the fine-tuning process. Therefore, it can help the model better adapt to the requirements of the news captioning task. First, we concatenate the rule vector $f$, the visual vector $v$, and the article vector $x^{A}$, which are input into the BART encoder:
$c=\mathrm{BART_{enc}}(f,v,x^{A})$ (8)
@@
Second, we inject the news-aware semantic rule vector $f$ into the last three encoder layers of BART (i.e., $M$ is set as 3 in Eq. 2). It is because compared with shallow encoder layers, deep ones (last three layers) are closer to the prediction outputs and have better semantic modeling capabilities. Therefore, embedding the news-aware semantic rule into the deep layers can boost the captioner to generate image descriptions following designated rule signal. It can be formulated as:
$\begin{split}X^{(l+1)}=&\mathcal{F}(\mathrm{LayerNorm}(\\ X^{(l)}+&\mathrm{MultiHead}([f;X^{(l)}],[f;X^{(l)}],[f;X^{(l)}])))\end{split}$ (9)
@@
where $X^{(l+1)}$ and $X^{(l)}$ are the input and output of the current encoder layer, respectively; $\mathcal{F}$ is the feed-forward layer; $\mathrm{LayerNorm}(\cdot)$ is the layer normalization; $f$ is the rule vector; [;][;][ ; ] represents the concatenation operation; $n$ is the number of parallel attention layers; $W_{i}^{Q}$, $W_{i}^{K}$, $W_{i}^{V}$, and $W^{O}$ are learnable weight matrices; $p$ is the prefix vector of the encoder.
@@ IV-A Experimental Setup
Training Details. We freeze the parameters of two large-scale pre-trained models, i.e., BART  [^3] https://github.com/openai/CLIP and CLIP(ViT-B/32)  [^4] https://huggingface.co/facebook/bart-large . The BART model comprises 12 encoder layers and 12 decoder layers, with each self-attention block incorporating 16 heads. The maximum length of input news text is set to 512 tokens, and each token has an embedding dimension of 1024. We assign a maximum length of 60 to the generated captions for the training and validation sets and 128 for the test set. Considering that the maximum input length of pre-trained named entity recognition model is 512 tokens, we perform slicing operations on the original news articles to ensure the model’s capability to handle the entire article. We employ the AdamW optimizer [47] and a linear learning rate scheduler during the training process. We set the batch size to 16, the training period to 30 epochs, and the learning rate to $5\times 10^{-5}$. The work is completed using the PyTorch framework, and the entire model is trained on two RTX 3090Ti GPUs. The training time per epoch is approximately 1.5 hours for the GoodNews dataset and approximately 3 hours for the NYTimes800k dataset.
@@ IV-B Comparison with the State-of-The-Art
  • Our method can significantly outperform all template-based methods and achieve competitive performance compared with end-to-end methods. Specifically, the template-based methods are limited by the capabilities of the template generator, which can not generate descriptions with diverse and expressive language. The end-to-end methods overlook the guidance of news-aware rule signal, resulting in the generated sentences that may not adhere to the fundamental rules of news reporting. Comparatively, our method builds the news-aware semantic rule to guide the model to generate news captions, which can accurately describe the individuals and actions associated with the event. As shown in Tab. I, on GoodNews, our method can outperform the best template-based competitor [5] by 59.82%, and the best end-to-end competitor [6] by 8.93% under CIDEr scores, respectively.
  • Tab. I shows our method can explicitly describe named entities, people’s names, and rare proper nouns against state-of-the-art methods. Thanks to the proposed news-aware semantic rule, which includes wide range of entities extracted from news articles, our method can achieve impressive results of 37.53% and 37.95% on GoodNews and NYTimes800k, respectively, under the recall of rare proper nouns. It makes the absolute improvement over the best competitor Tell [3] by 11.23% and 10.95%. However, we have observed that JoGANIC [4] and Zhou et al. [24] slightly outperform our method in terms of the precision score for named entities. It is because both of them benefits most from using the richer named entity information.

### 14 Mar 2024 v3

@@ Abstract
News Captioning, Rule-Driven, Large-Scale Pre-trained Model.
@@ I Introduction
To tackle this problem, we propose the rule-driven news captioning method, which can generate image descriptions following designated rule signal. Our method consists of three components: 1) Named Entity Extraction. We use the pre-trained named entity recognition (NER) technique to extract named entities from news articles. Then, the CLIP model is employed to compute the image-text similarity to match the named entities that closely align with the news image content. To ensure that the CLIP model is applicable to the news domain, we fine-tuned it on the news caption dataset. 2) News Rule Construction. We design the news-aware semantic rule for the descriptions. This rule incorporates the primary action depicted in the image (e.g., “performing”) and the set of roles played by named entities surrounding the action (e.g., “Agent” and “Place”). Meanwhile, it can present a clear affiliation between generic objects (e.g., “man”) in images and named entities (e.g., “Pedro Sa Moraes”) in news articles. 3) Caption Generation. We incorporate the news-aware semantic rule into the large-scale pre-trained model, BART, using the prefix-tuning strategy. Specifically, BART can be used as the pre-trained multi-modal summarization model [6] to generate captions for both images and news articles. We embed the semantic rule into BART’s multiple encoder layers. As a result, BART is effectively guided to generate news sentences that adhere to the rules of news reporting.
We conduct comprehensive experiments on two publicly available news datasets, i.e., GoodNews [5] and NYTimes800k [3]. Experimental results show that our method can achieve competitive or state-of-the-art performance against the existing methods on all benchmarks. Especially under CIDEr metric, our method can achieve significant improvement over the best competitor [4] by 7.59% on GoodNews. Ablation studies further examine the efficacy of each component. Besides, we visualize the generated news-aware semantic rules, to clearly observe the named entities’ changes improved by our method. In summary, we have the following main contributions:
@@ II-A General Image Captioning
The image captioning task is a challenging task that combines computer vision and natural language processing technology [7, 8, 9, 10]. Vinyals et al. [8] first use CNN to encode the image, and then use LSTM to generate the corresponding image description. Yao et al. [11] propose GCN [12], [13] plus LSTM. Li et al. [9] adopt a framework based on CNN-Transformer. They introduce EnTangled Attention in Transformer to integrate visual and semantic information more effectively. Compared with the CNN-Transformer framework, Liu et al. [14] achieve more efficient global context modeling through a full Transformer model. Zhang et al. [15] introduce the Adaptive-Attention module in the Transformer decoder to measure the contribution of visual cues and language cues to caption generation. Wu et al. [1] use segmentation features as additional information sources to enhance the contribution of visual information in the prediction process. Cornia et al. [16] design Universal Captioner, which adopts network-level noise data for training and uses the long-tail concept to describe images. Although traditional image captioning task performs very well, this method can only generate general descriptions and lacks specific contextual information [17, 10, 18]. In comparison, our method uses image-article pairs as input, and with the help of the background knowledge contained in the news articles, it can generate descriptions containing named entities, which enhances the practicality and accuracy of the descriptions.
@@ II-B News Image Captioning
In recent years, news image captioning task has received increasing attention [19], [3], [20]. Compared with the traditional image captioning task, the news image captioning task can effectively combine the image content and the background knowledge of the news article to generate captions containing named entities. Early research work often generates captions in two steps [5], [21]. Firstly, a caption with placeholders is generated, then the correct named entities are selected from news articles to replace the placeholders in the caption. For example, Biten et al. [5] use different types of named entities as placeholders to generate the template caption, e.g., PER performs in GPE on DATE. Then they select named entities from the sentences with the highest attention values in the news articles to fill these placeholders.
Recently, an increasing number of works have adopted end-to-end models. Tran et al. [3] utilize RoBERTa [22] and ResNet-152 [23] to encode news articles and images respectively, and handle uncommon words by using byte-pair encoding in the Transformer decoder. Zhao et al. [24] propose to mine external web knowledge to construct a multi-modal knowledge graph, thereby associating visual objects with named entities. Yang et al. [25] design a set of template components based on the six elements of news (i.e., who, when, where, what, why, how) to guide the model in generating news captions that comply with journalistic guidelines. Zhang et al. [4] utilize a multi-modal entity prompting mechanism to fine-tune the large-scale pre-trained model.
@@
Figure 2: Overview of the proposed method. (a) Named Entity Extraction. We use the Named Entity Recognition (NER) model to extract named entities and their categories from news articles. Then, the fine-tuned CLIP model is applied to calculate the image-text similarity score, filtering the named entities accordingly. (b) News Rule Construction. We design the news-aware semantic rule, which includes the primary action depicted in the image (e.g., “performing”) and the named entities involved in the action (e.g., “Ms. Micucci”, “Ms. Lindhome” and “theater”) along with their corresponding roles (e.g., “Agent” and “Stage”). (c) Caption Generation. We integrate the news-aware semantic rule, news article, and image, into BART, a large-scale pre-trained model, using the prefix-tuning strategy, for the news caption generation.
@@ II-C Large-Scale Pre-trained Models
Large-scale pre-trained models have received widespread attention [26], [27], [28], [29], [30], [31], [6]. These models are usually pre-trained on large-scale datasets, and then fine-tuned for specific tasks to better adapt to various complex application scenarios. They have demonstrated excellent results in tasks such as common sense reasoning [26], [27] and visual question answering [32], [33]. Among them, CLIP [31] leverages 400 million image-text pairs gathered from the internet, and uses contrastive learning [34], [35] to achieve cross-modal semantic alignment between image features and text features. BART [6] is a large-scale pre-trained language model based on Transformer. Our approach enables news description generation by injecting news-aware semantic rule into multiple encoder layers of BART, using the prefix-tuning strategy.
II-D Situation Recognition
Situation recognition aims to describe the main activities depicted in an image by using a semantic framework centered around verbs [36, 37, 38, 39] . This framework identifies the objects involved in the activities, as well as the roles these objects play within the activities. Cooray et al. [37] propose treating the prediction of semantic roles as a query-based visual reasoning problem. Pratt et al. [38] introduce the joint situation localizer, which aims to simultaneously predict semantic frames and groundings. Yu et al. [39] leverage diverse statistical knowledge to enable neural networks to perform adaptive global reasoning on nouns.
@@ III-A Problem Formulation
In order to generate captions that adhere to fundamental rules of news reporting, we first design the news-aware semantic rule $F$, which is filled using images and news articles. Then, we integrate this semantic rule into the large-scale pre-trained model (e.g., BART), where multiple encoder layers are embedded with news-aware semantic rule, to generate image descriptions. Formally, Eq. 1 is re-written as:
@@ III-B Named Entity Extraction
Named entities play a crucial role in enhancing the accuracy and richness of news captions by providing specific person’s name (PER), organization (ORG), and location (LOC). In this section, we extract named entities and their corresponding types from given news articles, which are then used in subsequent modules. We first employ a pre-trained named entity recognition (NER) model  [^1] https://huggingface.co/Davlan/distilbert-base-multilingual-cased-ner-hrl to extract a set of entities $E=\{e_{1},e_{2},\dots,e_{L}\}$ from the article $A$, where $L$ is the number of named entities. As shown in Fig. 2(a), to select the named entities that closely align with the news image content, we use the CLIP model [31] to evaluate the similarity between the image and each entity. However, due to significant discrepancies between the dataset used in the pre-training of CLIP and the data from the news domain, we fine-tune the CLIP model on two existing news caption datasets separately (i.e., GoodNews [5] and NYTimes800k [3]). Specifically, we define a correct image-text pair as a news image with its matching caption. Incorrect captions serve as negative samples. The CLIP model uses a contrastive loss function to increase the similarity of matching pairs and decrease it for mismatched pairs, enhancing the model’s skill in accurately linking images and text in news domain. The fine-tuned CLIP model is used to compute the similarity scores between the image $I$ and each entity $e_{i}\in E$, and then select the top-$k$ entities to build the new entity set $E_{k}$:
$E_{k}=\{e_{i}\}_{i=1}^{k}=\mathrm{arg~{}Top\mbox{-}k}(\mathrm{CLIP_{news}}(I,E))$ (3)
where $\mathrm{CLIP_{news}}(I,E)$ aims to calculate image-text matching scores by the fine-tuned CLIP model [31].
@@ III-D Caption Generation
We employ the frozen CLIP model to encode the image $I$ and the semantic rule $F$ independently, and obtain the visual features and the rule features, respectively. The former is embedded into the visual vector $z^{v}$ by a multi-layer perceptron (MLP), while the latter is transformed into the rule vector $z^{f}$ through a linear layer:
$z^{v}=\mathrm{MLP}(\mathrm{CLIP_{image}}(I))$ (6)
$z^{f}=\mathrm{Linear}(\mathrm{CLIP_{text}}(F))$ (7)
where $\mathrm{CLIP_{image}}(\cdot)$ is the image encoder; $\mathrm{CLIP_{text}}(\cdot)$ is the text encoder; For the article representation, we use the tokenizer to convert the news article into a collection of tokens $X^{A}=\{x_{1},x_{2},\dots,x_{N}\}$, where $x_{i}$ represents the $i$-th token in the article. We feed $X^{A}$ into an embedding layer and obtain the article vector $x^{A}$.
Essentially, the news captioning task requires adhering to some fundamental rules of news reporting, which is significantly different from the previous BART objective for the general summarization task. In this section, we utilize the prefix-tuning strategy [43]. This lightweight fine-tuning strategy involves optimizing a tiny task-specific vector called “prefix”, which can inject the task-specific structural information during the fine-tuning process. Therefore, it can help the model better adapt to the requirements of the news captioning task. First, we concatenate the rule vector $z^{f}$, the visual vector $z^{v}$, and the article vector $x^{A}$, which are input into the BART encoder:
$c=\mathrm{BART_{enc}}(z^{f},z^{v},x^{A})$ (8)
@@
Specially, we inject the news-aware semantic rule vector $z^{f}$ into the last three encoder layers of BART (i.e., $M$ is set as 3 in Eq. 2). It is because compared with shallow encoder layers, deep ones (last three layers) are closer to the prediction outputs and have better semantic modeling capabilities. Therefore, embedding the news-aware semantic rule into the deep layers can boost the captioner to generate image descriptions following designated rule signal. It can be formulated as:
$\begin{split}X^{(l+1)}=&\mathcal{F}(\mathrm{LayerNorm}(\\ X^{(l)}+&\mathrm{MultiHead}([z^{f};X^{(l)}],[z^{f};X^{(l)}],[z^{f};X^{(l)}])))\end{split}$ (9)
@@
where $X^{(l+1)}$ and $X^{(l)}$ are the input and output of the current encoder layer, respectively; $\mathcal{F}$ is the feed-forward layer; $\mathrm{LayerNorm}(\cdot)$ is the layer normalization; $z^{f}$ is the rule vector; [;][;][ ; ] represents the concatenation operation; $n$ is the number of parallel attention layers; $W_{i}^{Q}$, $W_{i}^{K}$, $W_{i}^{V}$, and $W^{O}$ are learnable weight matrices; $p$ is the prefix vector of the encoder.
@@ IV-A Experimental Setup
Training Details. The BART  [^3] https://huggingface.co/facebook/bart-large model comprises 12 encoder layers and 12 decoder layers, with each self-attention block incorporating 16 heads. The maximum length of input news text is set to 512 tokens, and each token has an embedding dimension of 1024. We assign a maximum length of 60 to the generated captions for the training and validation sets and 128 for the test set. In Eq. 3, k is set to 3. Considering that the maximum input length of pre-trained named entity recognition model is 512 tokens, we perform slicing operations on the original news articles to ensure the model’s capability to handle the entire article. We employ the AdamW optimizer [50] and a linear learning rate scheduler during the training process of BART. We set the batch size to 16, the training period to 30 epochs, and the learning rate to $5\times 10^{-5}$. The work is completed using the PyTorch framework, and the entire model is trained on two RTX 3090Ti GPUs. The training time per epoch is approximately 1.5 hours for the GoodNews dataset and approximately 3 hours for the NYTimes800k dataset. Besides, we fine-tune the CLIP using 421K image-caption pairs from GoodNews and 763K pairs from NYTimes800k. The learning rate is $5\times 10^{-6}$. The batch size is 50. The CLIP is fine-tuned on one RTX 3090Ti GPU for 2 hours on GoodNews and 3 hours on NYTimes800k.
@@ IV-B Comparison with the State-of-The-Art
  • Our method can significantly outperform all template-based methods and achieve competitive performance compared with end-to-end methods. Specifically, the template-based methods are limited by the capabilities of the template generator, which can not generate descriptions with diverse and expressive language. The end-to-end methods overlook the guidance of news-aware rule signal, resulting in the generated sentences that may not adhere to the fundamental rules of news reporting. Comparatively, our method builds the news-aware semantic rule to guide the model to generate news captions, which can accurately describe the individuals and actions associated with the event. As shown in Tab. I, on GoodNews, our method can outperform the best template-based competitor [5] by 58.48%, and the best end-to-end competitor [4] by 7.59% under CIDEr scores, respectively.
  • Tab. I shows our method can explicitly describe named entities, people’s names, and rare proper nouns against state-of-the-art methods. Thanks to the proposed news-aware semantic rule, which can adhere to the fundamental rules of news reporting, our method can achieve impressive results of 37.53% and 38.36% on GoodNews and NYTimes800k, respectively, under the recall of rare proper nouns. It makes the absolute improvement over the best competitor Tell [3] by 11.23% and 11.36%. However, we have observed that JoGANIC [25] and Zhou et al. [20] slightly outperform our method in terms of the precision score for named entities. It is because both of them benefit most from using the richer named entity information. In contrast, we focus solely on three types of named entities: person’s name (PER), organization (ORG), and location (LOC).